# Samba: Simple Hybrid State Space Models for Efficient Unlimited Context Language Modeling

### v1

@@ 3.2 Exploration on Attention and Linear Recurrence
Sliding RetNet 438M 24 4.32 10.38 9.96 9.87
@@
We pre-train all models on the same SlimPajama [SAKM+23] dataset under both around 438M and 1.3B settings, and evaluate these models by calculating perplexity on the validation set with context length at 4096, 8192, and 16384 tokens to investigate their zero-shot length extrapolation ability. Peak training throughput is also measured as an efficiency metric. The details of the hyperparameter settings are included in Appendix A. As shown in Table 3, Samba consistently outperforms all other models in different context lengths and model sizes. The training speed of Samba is competitive compared to pure Transformer-based models on the 1.3B scale. Mamba has significantly worse training throughput because Mamba layers have slower training speed than MLP layers, and the purebred Mamba models need to have more layers than other models at the same number of parameters. We can notice that the full attention-based model cannot extrapolate beyond its context length without specific length extrapolation techniques, which motivates us to use SWA for Samba. In Section 4, we further show that even hybridizing with one full attention layer will still lead to exploding perplexity at 16k sequence length. We can also find that while RetNet can extrapolate well under the 438M scale, it has an increasing perplexity on 16K length at the 1.4B scale, which may indicate that its input-independent decay may need specific tuning at different scales to work well.
@@ 3.3 Efficient Length Extrapolation
We use the test split of the Proof-Pile [ZAP22] dataset to evaluate the length extrapolation ability of our models at a scale of around 1.7B parameters. We follow Position Interpolation [CWCT23] for data pre-processing. The sliding window approach [PSL21] is used for the perplexity evaluation with a window size of 4096. Besides having the decoding throughput in Figure 1 for the generation efficiency metric, we also measure the prompt processing speed in Figure 3 for the models Samba 1.7B, Mistral 1.6B, Mamba 1.8B, Llama-3 1.6B and its Self-Extended [JHY+24] version SE-Llama-3 1.6B with the prompt length sweeping from 1K to 128K. We set the group size to 4 and the neighborhood window to 1024 for self-extension. We fix the total processing tokens per measurement to be 128K and varying the batch size accordingly. The throughput is measured on a single A100 GPU with the precision of bfloat16. We repeat the measurements 10 times and report the averaged results. We can see that Samba achieves 3.73×3.73$\times$3.73 × higher throughput in prompt processing compared to Llama-3 1.6B at the 128K prompt length, and the processing time remains linear with respect to the sequence length. We can also observe that the existing zero-shot length extrapolation technique introduces significant inference latency overhead on the full-attention counterpart, while it still cannot extrapolate infinitely with perplexity performance comparable to that of Samba. In Figure 1, we can also see that Mamba has a slowly and stably increasing perplexity up to 1M sequence length, which indicates that linear recurrent models can still not extrapolate infinitely if the context length is extremely large.
Beyond its efficiency in processing long context, Samba can also extrapolate its memory recall ability to 256K context length through supervised fine-tuning, and still keeps its linear computation complexity. We fine-tune Samba 1.7B on Passkey Retrieval with a 4K training sequence length for only 500 steps. As presented in Figure 4, Samba 1.7B demonstrates a remarkable ability to recall information from significantly longer contexts compared to Mistral 1.6B, a model based solely on Sliding Window Attention (SWA). This capability is particularly evident in the heatmap, where Samba maintains the perfect retrieval performance across a wider range of pass-key positions in a long document of up to 256K length. We also draw the training loss curve and the overall passkey retrieval accuracy across the fine-tuning procedure in Figure 6 and Figure 7 of Appendix B. We find that despite the fact that both architectures can reach near-zero training loss in less than 250 steps, Samba can achieve near-perfect retrieval early at 150 training steps, while the Mistral architecture struggles at around 30% accuracy throughout the training process. This shows that Samba can have better long-range retrieval ability than SWA due to the input selection mechanism introduced by the Mamba layers.
@@
3.4 Long-Context Understanding
The impressive results on the synthetic passkey retrieval task encourage us to perform full-cycle instruction tuning of the Samba-3.8B model. We follow the same post-training recipe used for the Phi-3-mini series and evaluate the downstream performance of the instruction-tuned Samba-3.8B-IT (preview) on both the long-context summarization tasks (GovReport [HCP+21], SQuALITY [WPC+22]) and the main short-context benchmarks (MMLU, GSM8K, HumanEval), as shown in Table 4. We can see that Samba has substantially better performance than Phi-3-mini-4k-instruct on both the short-context (MMLU, GSM8K, HumanEval) and long-context (GovReport) tasks, while still having the 2048 window size of its SWA layer and maintaining the linear complexity for efficient processing of long documents.
Table 4: Downstream performance comparison between instruction-tuned Samba 3.8B and Phi-3-mini-4K on both long-context and short-context tasks. We report 5-shot accuracy (averaged by category) for MMLU, 8-shot CoT [WWS+22] for GSM8K, 0-shot pass@1 for HumanEval, ROUGE-L for both GovReport and SQuALITY. † Results from the Phi-3 technical report [AJA+24].
Model MMLU GSM8K HumanEval GovReport SQuality
Phi-3-mini-4K-instruct † 68.8 82.5 58.5 14.4 21.6
Samba-3.8B-IT (preview) 71.9 87.6 62.8 18.9 21.2
@@ 4 Analysis
In this section, we analyze the experimental results of Samba by answering the following research questions. The perplexity results on SlimPajama have a fluctuation around ±$\pm 0.3\%$. Training speed is measured on 8$\times$×H100 GPUs by default. All the models in this section are trained on SlimPajama with 20B tokens and 4K sequence length, unless otherwise specified.
How to train models with Sliding Window Attention (SWA)?
Since SWA has linear complexity with respect to the sequence length, it seems alluring to trade off the batch size to have a longer training sequence length without substantially decreasing the training throughput. However, as shown in Table 5, when the sequence length is increased, the validation perplexity also increases in all context lengths due to smaller batch sizes, and the optimal ratio of sequence length/window size observed is 2, resulting in a training length of 4096.
Batch Size Sequence Length Training Speed Validation Context Length
(×$\times 10^{5}$ tokens/s) 2048 4096 8192 16384
1024 2048 (Full Attention) 10.4 11.59 38.12 156.18 357.32
512 4096 9.88 11.87 11.16 10.69 10.61
256 8192 9.66 11.98 11.26 10.79 10.69
128 16384 9.48 12.37 11.63 11.12 11.02
64 32768 9.29 12.94 12.46 11.96 11.86
@@
Some previous works [FDS+23, LLB+24] suggest a hybrid architecture of Mamba with full attention. However, as shown in Table 6, the extrapolation perplexity is exploding at a context length of 16k even if a single full attention layer is placed at the beginning of the model. Samba also has much better training throughput compared to Mamba-MLP alternatives because self-attention with the FlashAttention 2 implementation is more training efficient than Mamba when the sequence length is 4096.
@@
How many parameters should be allocated to Attention?
Given that Mamba can already capture low-rank information in the sequences through recurrent compression, the attention layers in Samba theoretically will only need to focus on information retrieval where a small number of attention heads should suffice. In Table 7, we explore the techniques of query head grouping [ALTdJ+23, Sha19], for both the Llama and Samba models. Surprisingly, both the Llama-2-SWA architecture and the Samba architecture show improved validation perplexity when there is only one key-value head. We conjecture that this is because small language models can be more easily optimized with fewer KV heads to pay attention to the contexts. We can also see that Samba has a 2×2$\times$2 × smaller optimal number of query heads than the SWA model, which confirms our hypothesis that Samba can support a smaller number of attention heads.
Table 7: Perplexity on SlimPajama of Llama-2-SWA and Samba models at the 430M scales trained with different number of Query and Key-Value heads. “KV Size” means the size of Key-Value vectors per token. Since grouped query attention will reduce the parameters for attention from $4d_{m}^{2}$ to roughly $2d_{m}^{2}$, we increase the hidden size of MLP from $8/3d_{m}$ to $46083d_{m}=$ 4608 to have roughly the same number of total parameters as the original models.
@@
Why hybrid is better?
We examine the entropy of the attention distributions for both the Samba 1.7B and the Mistral 1.6B models. As shown in Figure 5a, the Samba model has a larger variance of the attention entropy distributed over the layer indices, with an interesting pattern that the upper and lower layers have entropy higher than the middle layers. This may indicate that the attention layers are more specialized in the Samba architecture, with the middle layers focusing on precise retrieval with low-entropy attention, and the top and bottom layers focusing on integrating the global information through high-entropy attention. We can also see in Figure 5b that, compared to the Mamba-MLP model, Samba has a higher entropy of input selection probabilities in the middle layers. This indicates that, given the memory recalling ability of the attention layers, the Mamba layers can focus more on modeling the recurrent structure rather than performing retrieval with precise input selections. This kind of specialization can be beneficial for the downstream model performance, which may explain the impressive results from the Samba architecture. Details on how entropy is calculated are included in Appendix C.

### v2

@@ 3.2 Exploration on Hybridizing Attention and Linear Recurrence
Sliding RetNet 446M 24 4.32 10.38 9.96 9.87
Mega-S6 422M 24 3.26 12.63 12.25 12.25
Mamba-SWA-MLP 400M 24 4.21 10.07 9.67 9.59
MLP2-SWA-MLP 417M 24 5.08 10.95 10.50 10.41
Samba-NoPE 421M 24 4.48 10.11 28.97 314.78
@@
Since SSMs belong to a broader realm of linear recurrent models (Orvieto et al., 2023; Qin et al., 2023; Yang et al., 2023; Katsch, 2023; Qin et al., 2024), there exist multiple alternatives other than Mamba when combing attention-based layers with recurrent neural networks. We also add architecture ablation studies to justify the design choices of Samba. Specifically, in addition to Llama-2, Mamba, Samba and Mamba-SWA-MLP, we investigate the comparative analysis of the following architectures:
  • Llama-2-SWA is a pure attention-based architecture that replaces all full attention layers in Llama-2 with sliding window attention.
  • Sliding RetNet replaces Mamba layers in the Samba architecture with Multi-Scale Retention  (Sun et al., 2023) layers. RetNet is a linear attention model with fixed and input-independent decay applying to the recurrent hidden states.
  • Sliding GLA replaces Mamba layers in the Samba architecture with Gated Linear Attention (GLA) (Yang et al., 2023). GLA is a more expressive variant of linear attention with input-dependent gating.
  • Mega-S6 replaces all MD-EMA modules in the Mega (Ma et al., 2023) architecture with the ShortConv+S6 combinations from Mamba to adapt Mega to the modern Mamba architecture. Rotary position embedding, RMSNorm and Softmax attention are also adopted. We set the intermediate dimension of the Mega-S6 layer to be $d_{m}$ so that it has a roughly $5d_{m}^{2}$ number of parameters. This represents a classical baseline that conducts sequential intra-layer SSM-Attention hybridization.
  • MLP2-SWA-MLP replaces all Mamba layers in the Samba architecture to SwiGLU layers with $6d_{m}^{2}$ number of parameters.
  • Samba-NoPE removes the rotary relative position embedding in Samba and does not have any position embedding in the architecture.
@@
We pre-train all models on the same SlimPajama (Soboleva et al., 2023) dataset under both around 438M and 1.3B settings, and evaluate these models by calculating perplexity on the validation set with context length at 4096, 8192, and 16384 tokens to investigate their zero-shot length extrapolation ability. Peak training throughput is also measured as an efficiency metric. The details of the hyperparameter settings are included in Appendix G. As shown in Table 3, Samba consistently outperforms all other models in different context lengths and model sizes. The training speed of Samba is competitive compared to pure Transformer-based models on the 1.3B scale. Mamba has significantly worse training throughput because Mamba layers have slower training speed than MLP layers, and the purebred Mamba models need to have more layers than other models at the same number of parameters. Comparing Mamba-SWA-MLP with Samba, we can see that Samba has slightly better perplexity scores and higher training throughput. Mamba-SWA-MLP trades off the MLP layers with more I/O intensive Mamba and Attention layers, leading to slower training speed. This also indicates that Mamba-SWA-MLP will have slower decoding speed than Samba due to larger total cache size resulting from more SSMs and Attention layers. We can further observe that replacing Mamba with MLP speeds up the training but harms perplexity significantly, indicating the importance of Mamba layers in the Samba architecture. Interestingly, even though we use SWA in Samba architecture, Samba-NoPE still has exploded perplexities beyond its training length without RoPE. We can also find that while RetNet can extrapolate well under the 438M scale, it has an increasing perplexity on 16K length at the 1.4B scale, which may indicate that its input-independent decay may need specific tuning at different scales to work well.
Table 4: Downstream evaluation of models pre-trained with 100B tokens from SlimPajama. We measure the character-normalized accuracy for HellaSwag following Gu & Dao (2023). All tasks are evaluated in zero-shot.
Architecture Size ARC-Easy HellaSwag Wino. PIQA LAMBADA Avg.
acc $\uparrow$ acc_norm $\uparrow$ acc $\uparrow$ acc $\uparrow$ acc $\uparrow$
LLaMA-2 1.3B 55.09 52.32 53.35 71.11 48.52 56.08
LLaMA-2-SWA 1.3B 56.65 52.59 54.93 71.60 47.56 56.67
Sliding GLA 1.2B 56.94 52.52 56.75 71.38 48.17 57.15
Sliding RetNet 1.4B 57.66 52.64 56.75 71.33 48.34 57.34
Mega-S6 1.3B 50.63 41.91 52.96 68.17 37.88 50.31
Mamba 1.3B 58.08 54.93 53.99 71.98 45.97 56.99
Mamba-SWA-MLP 1.3B 59.64 54.50 55.25 72.42 49.12 58.19
MLP2-SWA-MLP 1.3B 55.18 50.32 52.80 70.67 48.11 55.42
Samba-NoPE 1.3B 58.38 54.62 56.51 72.03 51.08 58.52
Samba 1.3B 58.21 54.73 55.72 72.36 51.68 58.54
In Table 4, we evaluate all our 1.3B scale models on five typical commonsense reasoning tasks (ARC-Easy, HellaSwag, WinoGrande, PIQA and the OpenAI variant [^1] https://huggingface.co/datasets/EleutherAI/lambada_openai of LAMBADA (Paperno et al., 2016) ) to understand the effect of architecture designs on downstream performances. We can see that Samba has the best average accuracy, outperforming the LLaMA 2 architectures by a large margin. Similar to our perplexity evaluation, Samba and Samba-NoPE have similar average accuracies, whereas Mamba-SWA-MLP falls slightly behind. We observe that different architectures excel at different tasks. Mamba-SWA-MLP performs best on ARC-Easy, while Samba and Samba-NoPE achieve superior results on LAMBADA. Hybrid models based on Mamba generally outperform hybrid linear attention models and pure softmax-attention models on HellaSwag.
@@ 3.3 Efficient Length Extrapolation
(a) Perplexity on the test set of Proof-Pile
(b) Decoding throughput with batch size 16
Figure 2: Samba shows improved prediction up to 1M tokens in the Proof-Pile test set while achieving a 3.64$\times$ faster decoding throughput than the Llama-3 architecture on 64K generation length. We also include an SE-Llama-3 1.6B baseline which applies the SelfExtend (Jin et al., 2024) approach for zero-shot length extrapolation. All models are trained with 4K sequence length.
We use the test split of the Proof-Pile (Zhangir Azerbayev & Piotrowski, 2022) dataset to evaluate the length extrapolation ability of our models at a scale of around 1.7B parameters. We follow Position Interpolation (Chen et al., 2023a) for data pre-processing. The sliding window approach (Press et al., 2021) is used for the perplexity evaluation with a window size of 4096. Besides having the decoding throughput in Figure 2 for the generation efficiency metric, we also measure the prompt processing speed in Figure 6 of Appendix B for the models Samba 1.7B, Mistral 1.6B, Mamba 1.8B, Llama-3 1.6B and its Self-Extended (Jin et al., 2024) version SE-Llama-3 1.6B with the prompt length sweeping from 1K to 128K. We set the group size to 4 and the neighborhood window to 1024 for Self-Extension. We fix the total processing tokens per measurement to be 128K and varying the batch size accordingly. The throughput is measured on a single A100 GPU with the precision of bfloat16. We repeat the measurements 10 times and report the averaged results. We can see that Samba achieves 3.73×3.73$\times$3.73 × higher throughput in prompt processing compared to Llama-3 1.6B at the 128K prompt length, and the processing time remains linear with respect to the sequence length. We can also observe that the existing zero-shot length extrapolation technique introduces significant inference latency overhead on the full-attention counterpart, while it still cannot extrapolate infinitely with perplexity performance comparable to that of Samba. In Figure 2, we can also see that Mamba has a slowly and stably increasing perplexity up to 1M sequence length, which indicates that linear recurrent models can still not extrapolate infinitely if the context length is extremely large.
3.4 Long-Context Understanding
@@
Figure 4: Phonebook evaluation accuracy of different base models.
Beyond its efficiency in processing long context, Samba can also extrapolate its memory recall ability to 256K context length through supervised fine-tuning, and still keeps its linear computation complexity. We fine-tune Samba 1.7B on Passkey Retrieval with a 4K training sequence length for only 500 steps. As presented in Figure 4, Samba 1.7B demonstrates a remarkable ability to recall information from significantly longer contexts compared to Mistral 1.6B, a model based solely on Sliding Window Attention (SWA). This capability is particularly evident in the heatmap, where Samba maintains the perfect retrieval performance across a wider range of pass-key positions in a long document of up to 256K length. We also draw the training loss curve and the overall passkey retrieval accuracy across the fine-tuning procedure in Figure 7 and Figure 8 of Appendix C. We find that despite the fact that both architectures can reach near-zero training loss in less than 250 steps, Samba can achieve near-perfect retrieval early at 150 training steps, while the Mistral architecture struggles at around 30% accuracy throughout the training process. This shows that Samba can have better long-range retrieval ability than SWA due to the input selection mechanism introduced by the Mamba layers. In Figure 8, we can also notice that the pre-trained base Samba model has a retrieval accuracy (at step 0) similar to that of Mistral, highlighting the need for future work to improve Samba’s zero-shot retrieval capabilities.
The encouraging results on Passkey Retrieval drives us to further explore the limits of our finetuning approach. We perform instruction tuning to the Samba-3.8B base model on Phonebook (Jelassi et al., 2024) with only 100 steps on 4K sequence length and evaluate the resulting Samba-3.8B-FT model for a sequence length up to 8K. The evaluation setting requires the models to retrieve a random phone number from a phone book containing 20 (length 400) to 480 (length 8400) name-number pairs, resulting in a pressure test of memorization to Samba which has a constant memory state size. Surprisingly, as shown in Figure 4, we can see that the Samba-3.8B-FT model can close most of its gap with a full-attention model (Llama2 7B) that has twice the parameter size within the 4K training length, and achieves much better extrapolation accuracy compared to all other models including the Phi3 base model which also uses 2K sliding window attention. Since both Passkey Retrieval and Phonebook require models to remember numbers in a long context document, it is interesting to investigate if a model instruction-tuned on one task can transfer its ability to the other task in zero-shot. We directly evaluate the Passkey Retrieval finetuned Samba 1.7B and Mistral 1.6B models (named Samba 1.7B PK-FT and Mistral 1.6B PK-FT respectively) on the Phonebook task. As shown in Figure 4, Samba 1.7B has slightly better retrieval accuracy than Mistral 1.6B, but both models cannot generalize their number recall ability beyond its sliding window size. We leave it for future work to further explore the transferability of long-context capabilities in linear complexity models.
@@ 4 Analysis
In this section, we analyze the experimental results of Samba by answering the following research questions. The perplexity results on SlimPajama have a fluctuation around ±$\pm 0.3\%$. Training speed is measured on 8$\times$×H100 GPUs by default. All the models in this section are trained on SlimPajama with 20B tokens and 4K sequence length, unless otherwise specified. We also have additional analyses on the effectiveness of short convolution in Appendix D.
@@
Some previous works (Fu et al., 2023; Lieber et al., 2024) suggest a hybrid architecture of Mamba with full attention. However, as shown in Table 5, the extrapolation perplexity is exploding at a context length of 16K even if a single full attention layer is placed at the beginning of the model. Although hybridization with full attention in the second and middle sixth blocks (the fourth row in the table), following Dao et al. (2022b), can bridge the perplexity gap between full-attention hybrids and Samba, they still cannot extrapolate beyond the training sequence lengths. Samba also has much better training throughput compared to Mamba-MLP alternatives because self-attention with the FlashAttention 2 implementation is more training efficient than Mamba when the sequence length is 4096.
@@
Table 6: Perplexity on SlimPajama of Llama-2-SWA and Samba models at the 430M scales trained with different number of Query and Key-Value heads. “KV Size” means the size of Key-Value vectors per token and attention layer. Since grouped query attention will reduce the parameters for attention from $4d_{m}^{2}$ to roughly $2d_{m}^{2}$, we increase the intermediate size of MLP from $8/3d_{m}$ to $46083d_{m}=$ 4608 to have roughly the same number of total parameters as the original models.
@@
How many parameters should be allocated to Attention?
Given that Mamba can already capture low-rank information in the sequences through recurrent compression, the attention layers in Samba theoretically will only need to focus on information retrieval where a small number of attention heads should suffice. In Table 6, we explore the techniques of query head grouping (Ainslie et al., 2023; Shazeer, 2019), for both the Llama and Samba models. Surprisingly, both the Llama-2-SWA architecture and the Samba architecture show improved validation perplexity when there is only one key-value head. We conjecture that this is because small language models can be more easily optimized with fewer KV heads to pay attention to the contexts. We can also see that Samba has a 2×2$\times$2 × smaller optimal number of query heads than the SWA model, which confirms our hypothesis that Samba can support a smaller number of attention heads.
Potential explanations on why hybrid is better?
We examine the entropy of the attention distributions for both the Samba 1.7B and the Mistral 1.6B models. As shown in Figure 5a, the Samba model has a larger variance of the attention entropy distributed over the layer indices, with an interesting pattern that the upper and lower layers have entropy higher than the middle layers. This may indicate that the attention layers are more specialized in the Samba architecture, with the middle layers focusing on precise retrieval with low-entropy attention, and the top and bottom layers focusing on integrating the global information through high-entropy attention. We can also see in Figure 5b that, compared to the Mamba-MLP model, Samba has a higher entropy of input selection probabilities in the middle layers. This indicates that, given the memory recalling ability of the attention layers, the Mamba layers can focus more on modeling the recurrent structure rather than performing retrieval with precise input selections. This kind of specialization can be beneficial for the downstream model performance, which may explain the impressive results from the Samba architecture. Details on how entropy is calculated are included in Appendix E.
@@
5 Conclusion
Acknowledgement
Appendix A Related Works
Hybrid Recurrent Models
Many recent works (Park et al., 2024; Jelassi et al., 2024; Akyürek et al., 2024) point out the lack of retrieval ability of linear SSMs, and propose hybridization of SSMs with the Attention mechanism. However, the history of SSM/RNN-Attention hybridization can be directly dated back to the birth of the Attention mechanism (Bahdanau et al., 2014) which is proposed as a soft feature alignment technique for recurrent models to cope better with long sequences. The revitalization of the fact that linear recurrent models are sequentially parallelizable (Martin & Cundy, 2018; Gu et al., 2021) has catalyzed a contemporary renaissance in hybrid recurrent architectures. SPADE (Zuo et al., 2022), GSS (Mehta et al., 2023), MEGA (Ma et al., 2023), Block State transformers (Fathi et al., 2023) and Megalodon (Ma et al., 2024) combine SSMs with chunked attention, while H3 (Dao et al., 2022b), Mambaformer (Park et al., 2024) and Jamba (Lieber et al., 2024; Team et al., 2024) propose to hybridize with quadratic self-attention. Our works focus particularly on the wall-time efficiency and the length extrapolatability of the hybrid SSM-Attention models, and propose to interleave SSMs with Sliding Window Attention (SWA), which has both linear computation complexity and the translation-invariant property over the sequence length. Infini-Attention (Munkhdalai et al., 2024) is a recently proposed method that implements an intra-layer hybridization (Wu et al., 2022) between SWA and Linear Attention with the delta rule (Schlag et al., 2021). While the preliminary results look promising, its performance in the setting of large-scale pre-training from scratch remains questionable. The most similar work to ours is Griffin (De et al., 2024), which interleaves the Real-Gated Linear Recurrent Unit (RG-LRU) with Sliding Window Attention (SWA). However, Samba hybridizes SWA with Mamba instead of RG-LRU and shows that this simple hybrid architecture can provide substantially better performance over state-of-the-art Transformer architectures across scales, while Griffin and its follow-up work RecurrentGemma (Botev et al., 2024) only show comparable or worse results than Transformers. The original Mamba paper (Gu & Dao, 2023) also explores hybridizing pure Mamba models with full attention or MLP layers, but it does not consider the wall-time efficiency of these hybridization and only achieves marginally better performance than the pure Mamba model. In contrast, we are the first to show that interleaving Mamba with both SWA and MLP can substantially outperform modern Transformers (and Mamba) at a scale up to 3.8B parameters, while achieving comparable training speed and better length extrapolation ability under the perplexity metrics.
Efficient Sparse Attention
Previous works have proposed sparsifying self-attention (Vaswani et al., 2017) with a static attention pattern (Child et al., 2019; Zaheer et al., 2020; Beltagy et al., 2020) or a dynamic learnable pattern (Roy et al., 2020; Kitaev et al., 2020; Ren et al., 2023) to model long sequences with subquadratic complexity over the sequence length. However, due to the lack of hardware-aware efficient implementation, its actual wall-time training efficiency is often worse than the dense attention optimized with FlashAttention (Dao et al., 2022a; Dao, 2023; Shah et al., 2024). In this work, we choose Sliding Window Attention, a simple static sparse attention pattern, because it can easily leverage the highly optimized FlashAttention kernels to enjoy an actual training speed-up over its dense self-attention counterpart.
Length Extrapolation
Many previous works have focused on extending the context length of pretrained Transformers to improve their performance on long-context tasks. Methods such as LM-Infinite (Han et al., 2023), StreamingLLM (Xiao et al., 2024), and LongLoRA (Chen et al., 2023b) achieve linear complexity for length extrapolation, but they can only stabilize perplexity beyond the training sequence length rather than significantly improve it. In contrast, we demonstrate that pre-training Transformers with Sliding Window Attention from scratch enables natural improvements in perplexity beyond the training sequence length. Other approaches, including LLaMA-2-Long (Xiong et al., 2023), LongLLaMA (Tworkowski et al., 2023), PI (Chen et al., 2023a), LongRoPE (Ding et al., 2024) and Self-Extend (Jin et al., 2024), attempt to extend the full attention through modifying position embedding or continual training strategies, but they typically retain quadratic complexity in the attention mechanism with additional computation or memory I/O overhead, therefore they do not scale well to very long sequences. Although these methods achieve an improved perplexity on a sequence length that is multiple times longer than the training sequence length, their perplexity still explodes if the sequence is extremely long. Our method achieves both linear complexity and superior extrapolation performance compared to zero-shot length extrapolation methods, such as Self-Extend, under the perplexity metric. However, we acknowledge that, in terms of zero-shot retrieval performance, our method still lags behind these approaches. This underscores a trade-off between perplexity and retrieval performance in length extrapolation, which we plan to explore and address in future work.
Appendix B Additional Evaluation Results
In Table 7, we conduct comprehensive evaluations on a diverse subset of benchmarks to assess Samba 3.8B base model’s performance across all the domains mentioned in Section 3 to ensure a thorough examination of the model’s capabilities. We also report the performance of the Transformer++ (TFM++) model, which uses the same architecture, pre-training recipe as Phi3-mini, for a fair comparison. The details of the generation configurations are included in Appendix G. We compare with several strong baselines, including Llama 2 (Touvron et al., 2023), Mistral (Jiang et al., 2023), Mamba (Gu & Dao, 2023), Gemma (Team, 2024), Recurrent-Gemma (R-Gemma) (Botev et al., 2024), Llama 3 (MetaAI, 2024) and TFM++. As shown in Table 7, Samba achieves the highest average score on all benchmarks, demonstrating its superior performance in handling various language comprehension tasks. Notably, Samba excels in the GSM8K benchmark, achieving an absolute 18.1% higher accuracy than TFM++ trained on the same dataset. This shows the surprising complementary effect of combining SSM with the attention mechanism. We conjecture that when combined with attention, Mamba, as an input-dependent SSM, can focus more on performing the arithmetic operation through its recurrent states than on doing the retrieval operation which can be easily learned by the sliding window attention.
Table 7: Downstream performance comparison of the Samba 3.8B base model with other pretrained base language models without instruction tuning. ARC-C and HellaSwag are measured with character-normalized accuracy. MMLU and GSM8K are measured in 5-shot, while others are in zero-shot. We report the MC2 score for TruthfulQA, maj@1 for GSM8K, and pass@1 for HumanEval. Measured by ours. The fair comparison should only be considered between TFM++ and Samba.
Model Size Tokens MMLU Hella- ARC- Wino- Truth. GSM Hum. Avg.
Swag C Gran. QA 8K Eval
Llama 2 6.7B 2T 45.3 77.2 45.9 69.2 38.8 14.6 12.8 43.4
13B 2T 54.8 80.7 49.4 72.8 37.4 28.7 18.3 48.9
Mistral 7.2B - 60.1 81.3 55.5 75.3 42.2 35.4 30.5 53.6
Mamba 2.8B 600B 26.2 71.0 41.7 65.9 34.4 3.6 7.3 35.7
Gemma 2.5B 3T 42.3 71.4 42.1 65.4 33.1 17.7 22.0 42.0
8.5B 6T 64.3 81.2 53.2 72.3 44.8 46.4 32.3 56.4
R-Gemma 2.7B 2T 38.4 71.0 42.3 67.8 35.1 13.4 21.3 41.3
Llama 3 8.0B 15T+ 66.6 79.2 53.2 72.6 43.9 45.8 28.7 55.8
TFM++ 3.8B 3.2T 67.2 76.6 53.8 72.6 47.3 51.5 51.8 60.1
Samba 3.8B 3.2T 71.2 77.4 55.7 77.1 43.4 69.6 54.9 64.2
Table 8: Post-trained models quality on representative benchmarks under the chat mode. The fair comparison should only be considered between Samba and Phi3 as we control the training recipes and datasets to be the same. Best results are in bold, second best underlined.
Category Benchmark Samba (June) 3.8B Phi3 (June) 3.8B R-Gemma 9B FalconMamba 7B Jamba-1.5-Mini 12B/52B Llama-3.2-In 3B Llama-3.1-In 8B MMLU MMLU (5-shot) 69.0 67.2 60.5 62.1 69.7 61.8 68.1 MMLU-Pro (0-shot, CoT) 47.9 46.5 17.8 14.5 42.5 39.2 44 Reasoning ARC-C (10-shot) 87.8 86.8 52.0 62.0 85.7 76.1 83.1 GPQA (0-shot, CoT) 29.5 29.0 4.7 8.1 32.3 26.6 26.3 Math GSM8K (8-shot, CoT) 86.4 84.8 42.6 52.5 75.8 75.6 77.4 Code HumanEval (0-shot) 70.1 66.5 31.1 - 62.8 62.8 66.5 MBPP (3-shot) 71.7 70.0 42.0 - 75.8 67.2 69.4 Average 66.1 64.4 35.8 - 63.5 58.5 62.1
As shown in Table 8, we can see that post-trained hybrid models can achieve superior performance compared to industry-standard Transformer-based LLMs such as Llama-3.1-Instruct 8B and Llama-3.2-Instruct 3B, and SSM-based LLMs such as FalconMamba [^2] https://huggingface.co/tiiuae/falcon-mamba-7b-instruct. Recent progress on hybrid LLMs, including Jamba 1.5 (Team et al., 2024) and our own work on Samba, shows significant improvement over earlier approaches like R-Gemma (Botev et al., 2024), which hybridizes attention with linear recurrent models but is trained on smaller data scales. Samba delivers comparable performance to Jamba-1.5-Mini while using around 3$\times$ fewer active parameters and 13$\times$ fewer total parameters, due to an advanced text-book data synthesis technique (Abdin et al., 2024). Additionally, Samba outperforms the Phi3 architecture, which is trained on the same data and optimization setting, further highlighting the superiority of our hybrid architecture over modern Transformer models.
Appendix C Additional Experiment Details
We perform instruction tuning for both Mistral 1.6B and Samba 1.7B on Passkey Retrieval using document length 4096, where we generated the data on the fly through randomly sampling a 5-digit integer passkey value and a location/depth between zero and the document length to insert the passkey. The model is then asked to generate the passkey given the full document. We train both models using batch size 2048, 250 warm-up steps with a peak learning rate of $1e^{-4}$, and 0.1 weight decay with AdamW (Loshchilov & Hutter, 2018) optimizer. In both cases, the loss converges quickly in 100-200 steps. During the evaluation, we measure the overall average accuracies of the passkey retrieval at the document length of [4k, 8k, 16k, 32k, 64k, 128k, 256k], for each length we evaluate at 11 different depths of the document (from 0, 0.1, 0.2, … to 1.0). In addition, for each location of the passkey (depth) in the document, we evaluate the model with five different passkeys to measure accuracy. As seen in Figure 8, the average passkey retrieval accuracy for Samba 1.7B almost reaches 100% in around 150 steps, while the accuracy for Mistral 1.6B remains low, demonstrating the extrapolation ability of the Samba architecture.
Appendix D Additional Analyses
How to train models with Sliding Window Attention (SWA)?
Since SWA has linear complexity with respect to the sequence length, it seems alluring to trade off the batch size to have a longer training sequence length without substantially decreasing the training throughput. However, as shown in Table 9, when the sequence length is increased, the validation perplexity also increases in all context lengths due to smaller batch sizes (Varis & Bojar, 2021), and the optimal ratio of sequence length/window size observed is 2, resulting in a training length of 4096.
Batch Size Sequence Length Training Speed Validation Context Length
(×$\times 10^{5}$ tokens/s) 2048 4096 8192 16384
1024 2048 (Full Attention) 10.4 11.59 38.12 156.18 357.32
512 4096 9.88 11.87 11.16 10.69 10.61
256 8192 9.66 11.98 11.26 10.79 10.69
128 16384 9.48 12.37 11.63 11.12 11.02
64 32768 9.29 12.94 12.46 11.96 11.86